Circular No. 5689

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993a) Preliminary parabolic orbital elements from 18 observations, Jan. 2-6

(including means of the trail positions on IAUC 5687):

T = 1994 Jan. 2.303 TT  ω = 135.^{o}334
Ω = 144.413 }2000.0
q = 1.72810 AU  i = 124.668
1993TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 3 9^{h}46.^{m}09 +47^{o}11.^{′}9 3.731 4.512 138.^{o}2 8.^{o}4 14.4
8 9 39.84 +48 32.2 3.656 4.467 141.4 7.9 14.3
13 9 32.59 +49 52.0 3.589 4.422 143.9 7.5 14.2
18 9 24.33 +51 09.8 3.532 4.377 145.3 7.4 14.2
23 9 15.05 +52 23.9 3.484 4.331 145.5 7.4 14.1
28 9 04.83 +53 32.9 3.447 4.286 144.4 7.7 14.0
Feb. 2 8 53.79 +54 35.3 3.418 4.241 142.2 8.2 13.9
7 8 42.08 +55 29.9 3.400 4.195 139.1 8.9 13.9
12 8 29.92 +56 15.8 3.390 4.149 135.2 9.6 13.8

1H 0857–242 H. Barwig and W. Wimmer, Universitäts-Sternwarte, München; and

I. Bues, Remeis-Sternwarte, Bamberg, write: “Extensive high-speed photometry of the newly-detected cataclysmic variable 1H 0857–242 (IAUC 5481) has been performed with the European Southern Observatory 1-m telescope at La Silla and the MCCP fiberoptic multichannel photometer of the University of Munich. The star was in quiescence (V = 15.0 ± 0.5) during 1992 Dec. 26-1993 Jan. 1. It showed intense flickering activity at an amplitude in B of 40-60 percent of the mean brightness. On Jan. 2 UT, the object was found in outburst, having brightened 2.7 mag in B and 2.8 mag in V. Four CCD spectrograms obtained with the 1.52-m ESO telescope (+ Boller & Chivens spectrograph; range 360-700 nm, resolution 0.2 nm) show changes in flux and line structure. On Jan. 1.34, all the Balmer and He I lines appear in emission at intermediate strengths. The outburst spectrum of Jan. 2.34 reveals the Balmer lines with n ≥ 5 as well as He I (447.2 nm) in pure absorption, whereas Hβ has an emission core and Hα a P-Cyg profile. The two spectra of Jan. 3.25 and 3.35 demonstrate a steep decrease in intensity in the blue and a change from the absorption profiles via P-Cyg profiles to emission features stronger than those of Jan. 1.34.”

1993 January 7 Daniel W. E. Green